# Quantum Energy Teleportation under Equilibrium and Nonequilibrium Environments

### v1

@@ Abstract
Quantum energy teleportation (QET), implemented via local operations and classical communication, enables carrier-free energy transfer by exploiting quantum resources. While QET has been extensively studied theoretically and validated experimentally in various quantum platforms, enhancing energy output for mixed initial states, as the system inevitably interacts with environments, remains a significant challenge. In this work, we study QET performance in a two-qubit system coupled to equilibrium or nonequilibrium reservoirs. We derive an analytical expression for the energy output in terms of the system Hamiltonian eigenstates, enabling analysis of energy output for mixed states. Using the Redfield master equation, we systematically examine the effects of qubit detuning, nonequilibrium temperature difference, and nonequilibrium chemical potential difference on the energy output. We find that the energy output for mixed states often follows that of the eigenstate with the highest population, and that nonequilibrium environments can enhance the energy output in certain parameter regimes.
I Introduction
@@
Quantum teleportation (QT) is a well-known protocol that transmits the information of quantum states to remote locations using quantum entanglement together with local operations and classical communication (LOCC) [1, 2, 3]. Later, Hotta introduced a novel protocol called quantum energy teleportation (QET), which enables energy transmission via entanglement and LOCC [4]. Theoretically, QET can be realized in various physical systems, including spin chains [5, 6, 7], cold trapped ions [6], harmonic chains [8], and quantum fields [4, 9, 10, 11, 12]. QET has also been studied in holographic conformal field theory [13]. Recently, QET was experimentally demonstrated in the laboratory and on a quantum chip [14, 15]. Although QT and QET both rely on quantum correlations, their goals differ: QT transmits quantum state information, whereas QET aims to extract energy from a local subsystem rather than to restore the state.
In the original QET protocol [4], the sender (Alice) and the receiver (Bob) share the ground state, so Bob cannot extract energy from his subsystem by any local unitary operation alone. Alice’s measurement, however, increases the energy of her subsystem, and Bob’s subsystem collapses into a state that is no longer of minimal energy due to entanglement, enabling him to extract energy. The extracted energy is nevertheless less than the energy Alice injected, ensuring energy conservation. Alice must transmit her measurement outcome to Bob via classical communication before he can perform the extraction. Because there is no physical carrier transferring energy between Alice and Bob, the effective speed of energy transfer is limited only by the speed of classical communication.
@@
If Bob’s operations may include arbitrary local operations, he can extract more energy; this is referred to as strong QET [16]. Strong QET consistently yields greater extracted energy than standard QET, and they coincide only when Bob’s subsystem is left in a pure state after Alice’s measurement. Recently, Kazuki Ikeda proposed extending QET concepts beyond energy to arbitrary observables [17]. To illustrate this idea, he studied a (1+1)-dimensional Dirac system and used feedback control based on fermion chirality to activate electric current and charge, and he derived a rigorous upper bound on the teleported quantity. In conventional QET, the upper bound on energy output is severely constrained by distance; however, using squeezed vacuum states with local vacuum regions between the two parties can overcome this limitation [18]. In addition, a hyperbolic quantum network can realize long-range QET by transmitting local quantum information via quantum teleportation and performing conditional operations on that information [19].
@@
Energy teleportation necessarily requires quantum resources, but the specific resources relevant to QET depend on the setting and without an universal consensus. In the minimal QET model, Lin et al. found that initial-state entanglement and coherence show no clear relationship with the extractable energy, although they correlate positively with the energy-output efficiency [20]. Moreover, the change in system entropy during the measurement process sets a lower bound on the transferable energy [7]. For thermal states, QET is enabled by thermal discord [21]. However, in some cases quantum discord is not the resource for QET, as shown for a three-spin Ising chain in a Gibbs state [22]. Note that the total amount of transmitted energy and information is constrained by entanglement [23].
In practice, QET protocols inevitably involve environmental interactions. Since Alice and Bob are located separately, it is essential to account for the effects of distinct local environments on QET. In the standard QET model the system is assumed to be in the ground state, which yields both a low total energy transfer and low efficiency [24]. However, when the system is in a mixed state, the presence of excited populations need not be detrimental to energy extraction. Here we consider QET in a two-qubit model where each qubit interacts only with its own environment, and we investigate how equilibrium and nonequilibrium reservoirs can be exploited to improve QET performance.
@@
Notably, previous studies have shown that nonequilibrium environments can enhance various of quantum correlations, including the quantum entanglement [25, 26, 27, 28], quantum discord [29, 30, 31], quantum steering [32], Bell nonlocality [33], and temporal correlations [34, 35]. We consider QET under steady-state conditions, where, after completing a total protocol, the environment has “cooled” the system back to its initial state. Nonequilibrium steady states exhibit properties distinct from equilibrium cases [36, 37, 38, 39]. We apply the Bloch–Redfield master equation to describe the nonequilibrium two-qubit model, which enables us to simulate changes in energy output due to temperature or chemical potential differences between the baths [40, 41, 42, 43, 44, 45, 46]. Compared with the Lindblad master equation, the Bloch–Redfield equation, without the secular approximation, provides a more accurate description of nonequilibrium steady states [47, 48, 49, 50, 51, 52, 53]. The limitations of the Redfield equation regarding density-matrix positivity and methods to mitigate this issue are discussed in [43, 46, 54, 55]. Additionally, we consider the effect of detuning between the system energy levels, which can enhance nonequilibrium effects.
@@
We find that a temperature difference in bosonic reservoirs consistently suppresses QET, whereas in fermionic reservoirs a temperature difference can enhance QET. The chemical potential difference has a strong effect: when the average chemical potential is extreme (either much smaller or much larger than the system energy levels), QET is reduced; conversely, when the chemical potential is comparable to the system energy levels, QET can be enhanced within a certain range. For a system in a low-excitation state, increasing Alice’s energy level can improve the energy output, while for a system in a high-excitation state, increasing Bob’s energy level can likewise raise the energy output.
@@
The paper is organized as follows. In Sec. II we introduce the standard QET protocol and analyze the energy output when QET is performed on each eigenstate of the Hamiltonian. We also review the Redfield master equation used in our study. QET under equilibrium and nonequilibrium environments is analyzed in Secs. III and IV, respectively. Finally, in Sec. V we summarize our findings. For simplicity, we set $\hbar=k_{B}=1$ in the following sections.
II Energy Teleportation and Redfield Equation
@@
In this section, we first review the protocol of QET in Sec. II.1. Next, we analyze the energy output of the initial mixed state with an ”X” structure in Sec. II.2. Finally, we establish the model for our study in Sec. II.3, namely two qubits coupled to nonequilibrium environments.
II.1 Two-qubit Model of Energy Teleportation
@@
where $\varepsilon_{A,B}$ are the energy levels; $\kappa$ is interaction strength between the qubit $A$ and $B$; $\sigma^{z}_{A,B}$ and $\sigma^{x}_{A,B}$ are the Pauli operators of the qubits $A$ and $B$. In the original QET model, the energy level $\varepsilon_{A}$ equal to $\varepsilon_{B}$. We relax this constraint in our study, and consider detuning of the energy levels as an asymmetrical condition of system. Correspondingly the energy of the ground state is not necessarily zero [16]. The eigenvalues of the Hamiltonian in Eq. (1) are
@@
$E_{A}(\rho_{AB})=\sum_{u=\pm 1}\mathrm{Tr}\left(H_{AB}P_{A}(u)\rho_{AB}P_{A}^{\dagger}(u)\right).$ (8)
@@
$E_{B}(\rho_{AB})=\sum_{u=\pm 1}\mathrm{Tr}\left(H_{AB}U_{B}(u)P_{A}(u)\rho_{AB}P_{A}^{\dagger}(u)U_{B}^{\dagger}(u)\right).$ (10)
@@
II.2 Energy teleportation with $X$ state
@@
Consider the initial mixed state $\rho_{AB}$ with the “X”-structure, as expressed in the form
$\rho_{AB}^{X}=\begin{pmatrix}a&0&0&\alpha e^{i\beta}\\ 0&b&\delta e^{i\epsilon}&0\\ 0&\delta e^{-i\epsilon}&c&0\\ \alpha e^{-i\beta}&0&0&d\end{pmatrix}$, (13)
where all parameters are real and satisfy the normalization condition $a+b+c+d=1$ [58, 59]. Suppose that we adopt the Hamiltonian of system with the form in Eq. (1). Initially, prior to the QET protocol, the system in the state $\rho_{AB}^{X}$ has the energy
$E_{0}\left(\rho_{AB}^{X}\right)=(a+b-c-d)\varepsilon_{A}+(a-b+c-d)\varepsilon_{B}\\ +4\kappa(\delta\cos\epsilon+\alpha\cos\beta).$ (14)
@@
$E_{A}\left(\rho_{AB}^{X}\right)=(a-b+c-d)\varepsilon_{B}+4\kappa(\delta\cos\epsilon+\alpha\cos\beta),$ (15)
@@
$E_{B}\left(\rho_{AB}^{X}\right)\\ =\left((a-b+c-d)\varepsilon_{B}+4\kappa(\delta\cos\epsilon+\alpha\cos\beta)\right)\cos 2\theta\\ -2\left((-a+b-c+d)\kappa+\varepsilon_{B}(\delta\cos\epsilon+\alpha\cos\beta)\right)\sin 2\theta.$ (16)
@@
$D=2(-a+b-c+d)\kappa+\varepsilon_{B}(\delta\cos\epsilon+\alpha\cos\beta),$
$F=-(a-b+c-d)\varepsilon_{B}-4\kappa(\delta\cos\epsilon+\alpha\cos\beta).$
@@
II.3 Environments and Bloch-Redfield equation
@@
$H_{I}=\sum_{k_{A}}g_{k_{A}}\left(\sigma^{-}_{A}b_{k_{A}}^{\dagger}+\sigma^{+}_{A}b_{k_{A}}\right)\\ +\sum_{k_{B}}g_{k_{B}}\left(\sigma^{-}_{B}b_{k_{B}}^{\dagger}+\sigma^{+}_{B}b_{k_{B}}\right),$ (22)
@@
$H_{I}=\sum_{k_{A}}g_{k_{A}}(\eta_{A}+\xi_{A})b_{k_{A}}^{\dagger}\\ +\sum_{k_{B}}g_{k_{B}}(\eta_{B}+\xi_{B})b_{k_{B}}^{\dagger}+\text{H.c.},$ (23)
@@
$\frac{d\tilde{\rho}_{AB}}{dt}=-\int_{0}^{\infty}ds~\mathrm{Tr}_{R}\left[\tilde{H}_{I}(t),[\tilde{H}_{I}(t-s),\tilde{\rho}_{AB}\otimes\tilde{\rho}_{R}]\right],$ (26)
@@
The sign of $\beta_{j}(\varepsilon)$ is plus for the bosonic reservoirs, while it is minus for fermionic reservoirs. Parameters $T_{j}$ and $\mu_{j}$ are the equilibrium temperatures and chemical potentials of $j$-th reservoir, respectively. For bosonic reservoirs, such as photon or phonon baths, the particle number is not conserved with a vanishing chemical potential. Since the Bloch-Redfield equation is based on the assumption that the interaction between the system and the environment is weak, we can further assume that the coupling spectra with different frequencies are much less than the energy scale of the two qubits, namely $g_{k}\ll\varepsilon_{A},\varepsilon_{B}$. Therefore, it is reasonable to view $g_{k}$ as constants (independent of the transition frequencies $\varepsilon_{\pm}$).
The two-qubit steady state can be solved by reformulating the Bloch-Redfield equation in the Liouville space [28, 60]. It corresponds to the eigenstates of the superoperator with the zero eigenvalue. Because the four eigenstates of $H_{AB}$ is in $X$ form (13), the steady state either in the Hamiltonian eigenbasis or the local basis is an $X$ density operator.
III Quantum energy teleportation under equilibrium environments
We separately discuss the influence of bosonic and fermionic equilibrium environments on QET in Secs. III.1 and III.2 respectively.
III.1 Equilibrium bosonic environments
Figure 2: Energy output of steady states in the equilibrium bosonic reservoirs and the corresponding population of $|E_{1}\rangle$. ($\text{a}_{1}$) Energy output when the energy levels are set as $\varepsilon=0.1~(\text{black solid line})$, 1(red dashed line)1~($\text${red dashed line}), 5(green dot line)5~($\text${green dot line}) and 10(blue dashed dot line)10~($\text${blue dashed dot line}). ($\text{a}_{2}$) The population of $|E_{1}\rangle$ corresponding to ($\text{a}_{1}$). ($\text{b}_{1}$) Energy output when the temperatures are set as T=0.1(black solid line)T=0.1~($\text${black solid line}), 1(red dashed line)1~($\text${red dashed line}), 5(green dot line)5~($\text${green dot line}) and 10(blue dashed dot line)10~($\text${blue dashed dot line}). ($\text{b}_{2}$) The population of $|E_{1}\rangle$ corresponding to ($\text{b}_{2}$). The other parameters are set as $\kappa=1$ and $g_{A}=g_{B}=0.05$.

### v2

@@ Abstract
Quantum energy teleportation (QET), implemented via local operations and classical communication, enables carrier-free energy transfer by exploiting quantum resources. While QET has been extensively studied theoretically and validated experimentally in various quantum platforms, enhancing energy output for mixed initial states, as the system inevitably interacts with environments, remains a significant challenge. In this work, we study QET performance in a two-qubit system coupled to equilibrium or nonequilibrium reservoirs. Using the Redfield master equation, we systematically examine the effects of qubit detuning, nonequilibrium temperature difference, and nonequilibrium chemical potential difference on the energy output. We find that the energy output for mixed states often follows that of the eigenstate with the highest population, and that nonequilibrium environments can enhance the energy output in certain parameter regimes.
keywords:
Open quantum system , Quantum information protocol , Quantum energy teleportation
journal: Physics Letters A
1 Introduction
@@
Quantum teleportation (QT) is a well-known protocol that transmits the information of quantum states to remote locations using quantum entanglement together with local operations and classical communication (LOCC) [2, 5, 4]. Later, Hotta introduced a novel protocol called quantum energy teleportation (QET), which enables the extraction of energy from ground state via entanglement and LOCC [21]. Theoretically, QET can be realized in various physical systems, including spin chains [20, 22, 24], cold trapped ions [22], harmonic chains [39], and quantum fields [21, 25, 23, 11, 28]. QET has also been studied in holographic conformal field theory [12]. Recently, QET was experimentally demonstrated in the laboratory and on a quantum chip [45, 29]. Although QT and QET both rely on quantum correlations, their goals differ: QT transmits quantum state information, whereas QET aims to extract energy from a local subsystem rather than to restore the state.
In the original QET protocol [21], the sender (Alice) and the receiver (Bob) share the ground state, as a strong local passive (SLP) state [10, 1], Bob cannot extract energy from his subsystem by any local unitary operation alone. It is noteworthy that, despite the inclusion of ’teleportation’ in the protocol name, the protocol does not imply that the energy extracted by Bob directly originates from the energy injected by Alice. During the execution of the protocol, Alice’s measurement injects energy into her subsystem, leading to an increase only in her local energy. Simultaneously, Bob’s system undergoes a state change due to Alice’s measurement and can exhibit negative energy density. In this scenario, energy previously locked by the ground state is activated and can be extracted by Bob’s local operations. Over time, the energy injected by Alice’s measurement operations will infuse to Bob’s system through the interaction between the two subsystems. Crucially, in this model, Bob’s energy extraction occurs first, and only after this does the energy injected by Alice infuse to Bob via interaction. From Bob’s perspective, upon receiving Alice’s measurement results, he can extract energy from his system that was previously unable to yield it, as if Alice had transmitted energy to him. However, in reality, Alice merely activates the energy of Bob’s system through quantum entanglement. The activation of Bob’s energy, mediated by the collapse of the entangled state, is superluminal, while the actual energy extraction operation can only be performed after receiving Alice’s measurement result. In experimental verifications of quantum energy teleportation, the time taken for Bob to perform his operation is significantly shorter than the time it would take for energy to propagate to Bob’s system through interaction. This is a theoretical requirement and is also experimentally achievable. In this protocol, the energy extracted by Bob is less than the energy injected by Alice, thus not violating energy conservation.
@@
Generalized QET protocols have emerged prominently in recent research, such as, if Bob’s operations may include arbitrary local operations, he can extract more energy; this is referred to as strong QET [8]. Strong QET consistently yields extracted more energy from mixed states, and coincide with original QET protocol only when Bob’s subsystem is left in a pure state after Alice’s measurement. Furthermore, Kazuki Ikeda proposed extending QET concepts beyond energy to arbitrary observables [31]. To illustrate this idea, he studied a (1+1)-dimensional Dirac system and used feedback control based on fermion chirality to activate electric current and charge, and he derived a rigorous upper bound on the teleported quantity. In conventional QET, the upper bound on energy output is severely constrained by distance; however, using squeezed vacuum states with local vacuum regions between the two parties can overcome this limitation [19]. In addition, a hyperbolic quantum network can realize long-range QET by transmitting local quantum information via quantum teleportation and performing conditional operations on that information [30].
@@
QET necessarily requires quantum resources, but the specific resources relevant to QET depend on the setting and without an universal consensus. In the minimal QET model, Lin et al. found that initial-state entanglement and coherence show no clear relationship with the extractable energy, although they correlate positively with the energy-output efficiency [9]. Moreover, the change in system entropy during the measurement process sets a lower bound on the transferable energy [24]. For thermal states, QET is enabled by thermal discord [18]. However, in some cases quantum discord is not the resource for QET, as shown for a three-spin Ising chain in a Gibbs state [49]. Note that the total amount of transmitted energy and information is constrained by entanglement [50]. In this work, we analyze the behavior of QET from the perspective of energy eigenstates, rather than directly analyzing entanglement resources. This approach offers a more intuitive explanation for the variations in energy extraction, especially compared to the elusive entanglement resources.
@@
In practice, QET protocols inevitably involve environmental interactions. Since Alice and Bob are located separately, it is essential to account for the effects of distinct local environments on QET. In the standard QET model the system is assumed to be in the ground state, which yields both a low total energy transfer and low efficiency [16]. However, when the system is in a mixed state, the presence of excited populations need not be detrimental to energy extraction. Here we consider QET in a two-qubit model where each qubit interacts only with its own environment, and we investigate how equilibrium and nonequilibrium reservoirs can be exploited to improve QET performance. Notably, previous studies have shown that nonequilibrium environments can enhance various of quantum correlations, including the quantum entanglement [55, 6, 15, 53], quantum discord [38, 36, 42], quantum steering [57], Bell nonlocality [58], and temporal correlations [7, 59]. We consider QET under steady-state conditions, where, after completing a total protocol, the environment has “cooled” the system back to its initial state. Nonequilibrium steady states exhibit properties distinct from equilibrium cases [54, 34, 41, 46]. We apply the Bloch–Redfield master equation to describe the nonequilibrium two-qubit model, which enables us to simulate changes in energy output due to temperature or chemical potential differences between the baths [3, 44, 43, 32, 35, 40, 33]. Compared with the Lindblad master equation, the Bloch–Redfield equation, without the secular approximation, provides a more accurate description of nonequilibrium steady states [60, 37, 27, 61, 62, 52, 13]. The limitations of the Redfield equation regarding density-matrix positivity and methods to mitigate this issue are discussed in [32, 33, 47, 48]. Additionally, we consider the effect of detuning between the system energy levels, which can enhance nonequilibrium effects.
@@
We find that the temperature difference in bosonic reservoirs consistently suppresses QET, whereas in fermionic reservoirs the temperature difference can enhance QET. The chemical potential difference has a strong effect: when the average chemical potential is extreme (either much smaller or much larger than the system energy levels), QET is reduced; conversely, when the chemical potential is comparable to the system energy levels, QET can be enhanced within a certain range. For a system in a low-excitation state, increasing Alice’s energy level can improve the energy output, while for a system in a high-excitation state, increasing Bob’s energy level can likewise raise the energy output.
@@
The paper is organized as follows. In Sec. 2 we introduce the standard QET protocol and analyze the energy output when QET is performed on each eigenstate of the Hamiltonian. We also review the Redfield master equation used in our study. QET under equilibrium and nonequilibrium environments is analyzed in Secs. 3 and 4, respectively. Finally, in Sec. 5 we summarize our findings. For simplicity, we set $\hbar=k_{B}=1$ in the following sections.
2 Energy Teleportation and Redfield Equation
@@
In this section, we first review the protocol of QET in Sec. 2.1. Next, we analyze the energy output of the initial mixed state with an ”X” structure in Sec. 2.2. Finally, we establish the model for our study in Sec. 2.3, namely two qubits coupled to nonequilibrium environments.
2.1 Two-qubit Model of Energy Teleportation
@@
where $\varepsilon_{A,B}$ are the energy levels; $\kappa$ is interaction strength between the qubit $A$ and $B$; $\sigma^{z}_{A,B}$ and $\sigma^{x}_{A,B}$ are the Pauli operators of the qubits $A$ and $B$. The existence of interaction terms within the Hamiltonian does not necessarily entail that QET is a local phenomenon. It can be theoretically and experimentally demonstrated that QET preserves its non-local character, provided that Bob’s operational timeframe is shorter than the characteristic time associated with Alice’s energy infuse propagation timescale.
In the original QET model, the energy level $\varepsilon_{A}$ equal to $\varepsilon_{B}$. We relax this constraint in our study, and consider detuning of the energy levels as an asymmetrical condition of system. Correspondingly the energy of the ground state is not necessarily zero [8]. The eigenvalues of the Hamiltonian in Eq. (1) are
@@
$E_{A}(\rho_{AB})=\sum_{u=\pm 1}\text{Tr}\left(H_{AB}P_{A}(u)\rho_{AB}P_{A}^{\dagger}(u)\right).$ (8)
@@
$E_{B}(\rho_{AB})=\sum_{u=\pm 1}\text{Tr}\left(H_{AB}U_{B}(u)P_{A}(u)\rho_{AB}P_{A}^{\dagger}(u)U_{B}^{\dagger}(u)\right).$ (10)
@@
2.2 Energy teleportation with $X$ state
@@
Consider the initial mixed state $\rho_{AB}$ with the ’X’-structure, as expressed in the form
$\rho_{AB}^{X}=\begin{pmatrix}a&0&0&\chi e^{i\nu}\\ 0&b&\delta e^{i\epsilon}&0\\ 0&\delta e^{-i\epsilon}&c&0\\ \chi e^{-i\nu}&0&0&d\end{pmatrix}$, (13)
where all parameters are real and satisfy the conditions of density matrix [56, 17]. Suppose that we adopt the Hamiltonian of system with the form in Eq. (1). Initially, prior to the QET protocol, the system in the state $\rho_{AB}^{X}$ has the energy
$E_{0}\left(\rho_{AB}^{X}\right)=(a+b-c-d)\varepsilon_{A}+(a-b+c-d)\varepsilon_{B}+4\kappa(\delta\cos\epsilon+\chi\cos\nu).$ (14)
@@
$E_{A}\left(\rho_{AB}^{X}\right)=(a-b+c-d)\varepsilon_{B}+4\kappa(\delta\cos\epsilon+\chi\cos\nu),$ (15)
@@
$E_{B}\left(\rho_{AB}^{X}\right)\\ =\left((a-b+c-d)\varepsilon_{B}+4\kappa(\delta\cos\epsilon+\chi\cos\nu)\right)\cos 2\theta\\ -2\left((-a+b-c+d)\kappa+\varepsilon_{B}(\delta\cos\epsilon+\chi\cos\nu)\right)\sin 2\theta.$ (16)
@@
$D=2(-a+b-c+d)\kappa+\varepsilon_{B}(\delta\cos\epsilon+\chi\cos\nu),$
$F=-(a-b+c-d)\varepsilon_{B}-4\kappa(\delta\cos\epsilon+\chi\cos\nu).$
@@
2.3 Environments and Bloch-Redfield equation
@@
$H_{I}=\sum_{k_{A}}g_{k_{A}}\left(\sigma^{-}_{A}b_{k_{A}}^{\dagger}+\sigma^{+}_{A}b_{k_{A}}\right)+\sum_{k_{B}}g_{k_{B}}\left(\sigma^{-}_{B}b_{k_{B}}^{\dagger}+\sigma^{+}_{B}b_{k_{B}}\right),$ (22)
@@
$H_{I}=\sum_{k_{A}}g_{k_{A}}(\eta_{A}+\xi_{A})b_{k_{A}}^{\dagger}+\sum_{k_{B}}g_{k_{B}}(\eta_{B}+\xi_{B})b_{k_{B}}^{\dagger}+\text{H.c.},$ (23)
@@
$\frac{d\tilde{\rho}_{AB}}{dt}=-\int_{0}^{\infty}ds~\text{Tr}_{R}\left[\tilde{H}_{I}(t),[\tilde{H}_{I}(t-s),\tilde{\rho}_{AB}\otimes\tilde{\rho}_{R}]\right],$ (26)
@@
The sign of $\beta_{j}(\varepsilon)$ is plus for the bosonic reservoirs, while it is minus for fermionic reservoirs. Parameters $T_{j}$ and $\mu_{j}$ are the equilibrium temperatures and chemical potentials of $j$-th reservoir, respectively. For bosonic reservoirs, such as photon or phonon baths, the particle number is not conserved with a vanishing chemical potential. Since the Bloch-Redfield equation is based on the assumption that the interaction between the system and the environment is weak, we can further assume that the coupling spectra with different frequencies are much less than the energy scale of the two qubits, namely $g_{k_{j}}\ll\varepsilon_{A},\varepsilon_{B}$. Therefore, it is reasonable to view $g_{k_{j}}$ as constants (independent of the transition frequencies $\varepsilon_{\pm}$), we set $g_{k_{A}}=g_{A},g_{k_{B}}=g_{B}$.
The two-qubit steady state can be solved by reformulating the Bloch-Redfield equation in the Liouville space [53, 51]. It corresponds to the eigenstates of the superoperator with the zero eigenvalue. Because the four eigenstates of $H_{AB}$ is in $X$ form (13), the steady state either in the Hamiltonian eigenstates or the local basis is an $X$ density operator.
3 Quantum energy teleportation under equilibrium environments
We separately discuss the influence of bosonic and fermionic equilibrium environments on QET in Secs. 3.1 and 3.2 respectively.
3.1 Equilibrium bosonic environments
Figure 2: Energy output of steady states in the equilibrium bosonic reservoirs and the corresponding population of $|E_{1}\rangle$. ($\text{a}_{1}$) Energy output when the energy levels are set as $\varepsilon=0.5~(\text{black solid line})$, 1(red dashed line)1~($\text${red dashed line}), 2(green dot line)2~($\text${green dot line}) and 5(blue dashed dot line)5~($\text${blue dashed dot line}). ($\text{a}_{2}$) The population of $|E_{1}\rangle$ corresponding to ($\text{a}_{1}$). ($\text{b}_{1}$) Energy output when the temperatures are set as T=0.1(black solid line)T=0.1~($\text${black solid line}), 1(red dashed line)1~($\text${red dashed line}), 5(green dot line)5~($\text${green dot line}) and 10(blue dashed dot line)10~($\text${blue dashed dot line}). ($\text{b}_{2}$) The population of $|E_{1}\rangle$ corresponding to ($\text{b}_{2}$). The other parameters are set as $\kappa=1$ and $g_{A}=g_{B}=0.05$.